3x+8y=4
3/2x+4y=2 i need the answer

Answer by stanbon(75887)

1st: 3x+8y=4
2nd: 3/2x+4y=2
-------------------
Multiply thru the 2nd equation by 2 to get:
3rd: 3x + 8y = 4
------------------------
This shows that the 2nd equation is a multiple of the 1st equation.
The system of equations is said to be dependent.
The solution for the system is all the points in the line 3x+8y=4
or [x , (-3/8)x+(1/2)]
==================
Cheers,
Stan H.